OpenStudy (anonymous):

Help please? Determine the number of cycles each trigonometric function has in the interval from 0 to 2π. Find the amplitude and period of each function. a. y = sin5θ b. y = -5sin2πθ c. y = 2cos4θ d. y = 4cos2θ

OpenStudy (campbell_st):

in general sin and cos can be written \[y = acos(bx)...... or ..... y = asin(bx)\] a is the amplitude of the function. e.g. y = 3sin(x) has an amplitude of 3 y = -2cos(x) has an amplitude of 2 and the curve has been reversed. the period is found using \[period = \frac{2\pi}{\left| b \right|}\] so for y = cos(5x) the period or length of the cycle is \[\frac{2\pi}{5}\] hope this helps